# PriorDiffusion: Leverage Language Prior in Diffusion Models

### v1

@@ 4 Experiments
$\tilde{y*}=\left(\frac{y^{*}-y_{2}}{y_{2}-y_{98}}-0.5\right)\times 2,$ (1)
@@
Evaluation datasets. We assess the performance of our PriorDiffusion using 4 real-world datasets that were not part of its training data, following the configuration as Marigold [35]. Both NYUv2 [69] and ScanNet [10] consist of indoor scenes captured with an RGB-D Kinect sensor. For NYUv2, we use the specified test set, containing 654 images. With ScanNet, we randomly select 800 images from the 312 official validation scenes for evaluation. The KITTI dataset [24, 25] features outdoor street scenes with sparse metric depth captured by a LiDAR sensor, where we apply the Eigen test split [14] consisting of 652 images, following the evaluation protocol of [15]. For ETH3D [68], which are also based on LiDAR measurements, we use all 454 samples that include ground truth depth maps.
Evaluation metric. Following the affine-invariant depth evaluation protocol [97, 60, 88, 35, 59], for each image and the predicted relative depth $y$, we fit a pair of scalars denoting the scale and shift parameters of the transformation: $(\hat{\alpha},\hat{\beta})=g_{\psi}(y,y^{*})\in\mathbb{R}^{2}$. The metric depth prediction is obtained by $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ such that:
and mean absolute relative error, calculated as:
Figure 5: Convergence speed comparison. Our method converges faster during training compared with Marigold, since it’s more efficient to learn 3D property from condensed, low-dimensional language prior. Due to computation overhead, following the same configuration of Marigold, we use a subset of KITTI and NYUv2 for training-time evaluation.
@@
Implementation details. We implemented our method using PyTorch, employing Stable Diffusion v2 [63] as the backbone and adhering to the original pre-training setup with the $v$-objective [65]. During training, we used the DDPM noise scheduler [30] with 1000 diffusion steps, while at inference time, the DDIM scheduler [71] was applied with 50 sampling steps. Our training process spanned 30000 iterations with a batch size of 32, achieved through gradient accumulation over 16 steps (batch size of 2 per step) to fit a single GPU. We optimized the model using the Adam optimizer with a learning rate of $3\cdot 10^{-5}$ and included random horizontal flipping as data augmentation. Training to convergence (usually our method converges after 20000 iterations) required approximately 4 days on an Nvidia RTX 3090 GPU.
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Blank text input 95.5 6.2 89.3 10.9 95.0 6.9 “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 89.9 10.7 95.3 6.8 Template B 95.7 6.1 89.8 10.7 95.2 6.8 Template C 95.8 6.1 89.8 10.7 95.3 6.8 Ours 95.9 5.9 90.6 10.4 95.7 6.5
  • a
  • b
  • c
Table 2: Template prompt comparison. Metrics are reported as percentages. PriorDiffusion can preserve performance when language input is not feasible thus using a fixed prompt instead.
@@
Quantitative comparison. In Table 1, we compare our PriorDiffusion against other state-of-the-art affine-invariant depth estimators across all those zero-shot benchmarks listed in the previous section. PriorDiffusion achieves the highest $\delta_{1}$ accuracy and the lowest absolute relative error (AbsRel) on all tested datasets, including NYUv2, KITTI, ETH3D, and ScanNet. Notably, PriorDiffusion matches or surpasses top competitors such as Marigold and DPT, demonstrating strong generalization capabilities with an AbsRel of 5.9% and $\delta_{1}$ of 95.9% on NYUv2, and comparable excellence across other benchmarks. This is attributed to the generalization ability of language due to its invariance and robustness across diverse 3D scenes. As shown in Figure 5, another significant advantage of PriorDiffusion is its faster convergence, compared to Marigold. In Table 1, even at 20,000 iterations, PriorDiffusion achieves performance metrics comparable to or better than Marigold’s results at 30,000 iterations. This can be attributed to the language prior integrated into our model, which provides additional semantic and geometric prior serving as extra constrain that speed up the convergence diffusion process. An intuitive explanation is that the model expends less effort in learning the geometric properties and spatial relationships of objects from the image, which is high-dimensional and computationally intensive to process. Instead, much of this feature is already embedded in the language input, where these features are represented in a condensed and low-dimensional format. Human-provided language serves as an effective prior, allowing diffusion trajectory to converge more quickly.
@@
Template prompt comparison. In situations where user-provided descriptions are not available, we aim to explore whether our method can still perform effectively using either a blank input or standardized template prompts. If no specific language description that aligns with the scene is given by the user, we could instead use some pre-defined template prompt, such as “An image” or more elaborate prompts like, “A complex 3D scene with varying objects at different distances.”, as input of PriorDiffusion to maintain its performance. ” These templates, as presented in Table 2, assist in preserving the model’s performance when language input is not practical. It is important to note that we did not account for meaningless language inputs during training and inference. Consequently, using template prompts as input during inference may differ from the training domain, leading to a significant covariate shift in language descriptions. This issue can be mitigated by incorporating template prompts as part of the input during training, enabling PriorDiffusion to better adapt to various scenarios where meaningful language descriptions are not feasible.
@@ 5 Discussion
Limitation. One primary limitation is its dependence on the accuracy and detail of the provided language descriptions. Ambiguous or misleading text inputs from human users can result in suboptimal depth predictions and potentially compromise the model’s performance. Additionally, the computational cost of training and inferring using diffusion models remains higher than that of other CNN or Transformer-based methods, which may affect the feasibility of deploying this approach in real-time or resource-constrained environments. Last, our method requires language input from users, which may not always be feasible. In cases where user-provided descriptions are unavailable, our method can still function standardized template prompt. Another potential solution is to generate language descriptions on the fly using a vision-language model [11], an image captioner [31], or by following methods like RSA [97] to create structured text from the output of vision models [115, 39]. While we have not explored this approach in this paper, it presents a promising avenue for future research.
Conclusion. We present PriorDiffusion, a novel approach that integrates language priors with diffusion models to enhance monocular depth estimation. By leveraging geometric prior learned by text-to-image model that associated with language, PriorDiffusion overcomes traditional ambiguities and visual nuisances associated with monocular depth estimation, achieving state-of-the-art performance across various zero-shot benchmarks, while maintaining a faster convergence speed. We also highlights the potential of bridging the gap between visual perception and language understanding in vision-based AI systems.
@@ B.1. Visualization
We implemented our method using PyTorch, employing Stable Diffusion v2 [63] as the backbone and maintaining the original pre-training configuration with the $v$-objective [65]. The training process utilized the DDPM noise scheduler [30] with 1,000 diffusion steps, while at inference time, the DDIM scheduler [71] was employed with 50 sampling steps for faster results. Our training setup spanned 30,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (with a per-step batch size of 2) to fit on a single Nvidia RTX 3090 GPU. We used the Adam optimizer with a learning rate set at $3\cdot 10^{-5}$ and included random horizontal flipping with a probability of 0.5 as data augmentation. For depth normalization, we employed a scale and shift-invariant method with clipping enabled and set the normalization range between -1.0 and 1.0, using a 0.02 min-max quantile to maintain robustness. This normalization strategy was applied during training and zero-shot evaluations for consistency. The training noise scheduler initialized from the pre-trained Stable Diffusion v2 [63] model maintained a noise strength of 0.9 and incorporated an annealed strategy to progressively reduce noise levels. We saved checkpoints every 50 iterations, with backup, validation, and visualization checkpoints set at intervals of 2,000 iterations. The training process typically converged after approximately 20,000 iterations, though we extended training to 30,000 iterations for thorough coverage. We used mean squared error (MSE) as the loss function, with reduction set to “mean” for averaged loss calculation. A customized iteration-wise exponential scheduler is applied which adjusts the learning rate iteratively using an exponential decay function. It decays the learning rate to 1% of its initial value over 25,000 iterations with a warmup phase of 100 steps.
@@ C.2. Ablation for prompts to generate text
To mimic the language description that humans might provide to assist depth estimator during practical usage, we use the visual question-answering model LLaVA v1.6 Mistral [47]. Here we want to study the effect of different prompts. The prompt we use should elicit responses that capture essential details, including the positioning of objects, their interactions, and notable features that influence monocular depth estimation. We generate different prompts using ChatGPT 4o [52], with the prompt:

### v2

@@ 4 Experiments
$\tilde{y*}=\left(\frac{y^{*}-y_{2}}{y_{98}-y_{2}}-0.5\right)\times 2,$ (1)
@@
Evaluation datasets. We assess the performance of our PriorDiffusion using 4 real-world datasets that were not part of its training data, following the configuration as Marigold [38]. Both NYUv2 [72] and ScanNet [10] consist of indoor scenes captured with an RGB-D Kinect sensor. For NYUv2, we use the specified test set, containing 654 images. With ScanNet, we randomly select 800 images from the 312 official validation scenes for evaluation. The KITTI dataset [25, 26] features outdoor street scenes with sparse metric depth captured by a LiDAR sensor, where we apply the Eigen test split [14] consisting of 652 images, following the evaluation protocol of [15]. For ETH3D [71], which are also based on LiDAR measurements, we use all 454 samples that include ground truth depth maps. Following the affine-invariant depth evaluation protocol [99, 63, 90, 38, 62], we evaluate first-order threshold accuracy $\delta_{1}$ and mean absolute relative error $AbsRel$. Details of the metric are provided in the supplementary material.
Obtain language description. To implement our method, we require human-provided text descriptions for each image. Since standard benchmarks do not include such descriptions, we propose using off-the-shelf models to generate text descriptions to simulate those a human would provide. For this purpose, we use the visual question-answering model LLaVA v1.6 [49] to generate 10 text descriptions for each training image using different prompts, and 1 text description for each testing image for evaluation. The details of prompt design are shown in the supplementary material. In each training step, we randomly select 1 text out of 10 that aligns with the input image.
@@
Implementation details. We implemented our method using PyTorch, employing Stable Diffusion v2 [66] as the backbone and adhering to the original pre-training setup with the $v$-objective [68]. During training, we used the DDPM noise scheduler [33] with 1000 diffusion steps, while at inference time, the DDIM scheduler [74] was applied with 50 sampling steps. Our training process spanned 30000 iterations with a batch size of 32, achieved through gradient accumulation over 16 steps (batch size of 2 per step) to fit a single GPU. We optimized the model using the Adam optimizer with a learning rate of $3\cdot 10^{-5}$ and included random horizontal flipping as data augmentation.
@@
Quantitative comparison. Shown in Table 3, PriorDiffusion matches the performance of state-of-the-art affine-invariant depth estimators that require significantly more training images, while outperforming methods trained with a similar number of images. Another significant advantage of PriorDiffusion is its faster convergence. Compared to Marigold, even at 20,000 iterations, PriorDiffusion achieves performance metrics comparable to or better than Marigold’s results at 30,000 iterations. This is also demonstrated in Figure 7. This can be attributed to the language prior integrated into our model, which provides additional semantic and geometric prior serving as an extra constrain that speed up diffusion training. An intuitive explanation is that the model expends less effort in learning the geometric properties and spatial relationships of objects from the image, which is high-dimensional and computationally intensive to process. Instead, much of these features are already embedded in the language input, where these features are represented in a condensed and low-dimensional format, which is easier to learn.
@@
Faster training convergence. One of the key advantages of PriorDiffusion is its faster convergence compared to Marigold. As shown in Figure 7 here, the improved convergence can be attributed to the integration of a language prior in our model, which provides additional semantic and geometric constraints that help accelerate the diffusion process. Due to computation overhead, following the same configuration of Marigold, we use a subset of KITTI and NYUv2 for training-time evaluation.
Figure 7: Convergence speed comparison. PriorDiffusion converges faster during training compared with Marigold, since it’s more efficient to learn 3D property from condensed, low-dimensional language feature.
Figure 8: Performance under different denoising steps. PriorDiffusion consistently outperforms Marigold across various denoising steps, with a faster convergence.
Fewer denoising steps. As demonstrated in Figure 8, and detailed in Table 4 in the supplementary material, we evaluate PriorDiffusion and Marigold with different denoising steps during inference. PriorDiffusion consistently outperforms Marigold across different denoising steps. Remarkably, PriorDiffusion converges in just 10 steps, while Marigold takes 25 steps to achieve convergence. This indicates that the language prior serves as an effective initialization of the diffusion trajectory, enabling faster convergence.
Better perception in small areas. To demonstrate a better perception for small objects and areas mentioned in text, as shown in Figure 4. We use MaskDINO to obtain panoptic segmentation masks for all images in NYUv2 test set, then evaluate the performance for masks that occupied less than 5%, 10%, and 20% of an image. Table 2 shows that small areas are harder to perceive compared with large areas, and PriorDiffusion obtains better improvement over small areas, as language prior provides geometric and semantic properties that guide the diffusion model to better perceive small areas and tiny structures.
NYUv2 Standard Small Area (5%) Small Area (10%) Small Area (20%) $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 95.7 6.1 91.7 9.0 92.2 8.4 93.9 7.3 Ours 95.9 5.9 92.8 8.3 93.1 7.9 94.6 6.8
Table 2: Better perception in small areas. We evaluate the performance for panoptic segmentation masks that occupied less than 5%, 10%, and 20% of an image. PriorDiffusion obtains better improvement over small areas due to geometric properties of those areas offered by the language prior.
Figure 9: Failure cases. Incorrect text (here we replace ”a bookshelf with glass” with ”a window with curtains”) misguides PriorDiffusion to perceive the geometry of objects.
@@ 5 Discussion
Conclusion. We present PriorDiffusion, a novel approach that integrates language prior with diffusion models to enhance monocular depth estimation. By leveraging geometric prior learned by text-to-image model that associates with language, PriorDiffusion overcomes traditional ambiguities and visual nuisances associated with monocular depth estimation, achieving state-of-the-art performance across various zero-shot benchmarks, while maintaining a faster convergence for training and inference diffusion trajectory.
Limitation. One primary limitation is its dependence on the accuracy and detail of the provided language descriptions. Shown in Figure 9, ambiguous or misleading text inputs from human users can result in suboptimal depth predictions and potentially compromise the model’s performance. Last, our method requires language input from users, which may not always be feasible. In cases where user-provided descriptions are unavailable, as demonstrated in Figure 5 in the supplementary material, our method can still function with standardized template prompts. Another potential solution is to generate language descriptions on the fly using a vision-language model [11], an image captioner [34], or by following methods like RSA [99] to create structured text from the output of visual algorithm [116, 42]. While we have not explored this approach in this paper, it presents a promising avenue for future research.
@@ B.1. Visualization
We implemented our method using PyTorch, employing Stable Diffusion v2 [66] as the backbone and maintaining the original pre-training configuration with the $v$-objective [68]. The training process utilized the DDPM noise scheduler [33] with 1,000 diffusion steps, while at inference time, the DDIM scheduler [74] was employed with 50 sampling steps for faster results. Our training setup spanned 30,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (with a per-step batch size of 2) to fit on a single Nvidia RTX 3090 GPU. We used the Adam optimizer with a learning rate set at $3\cdot 10^{-5}$ and included random horizontal flipping with a probability of 0.5 as data augmentation. For depth normalization, we employed a scale and shift-invariant method with clipping enabled and set the normalization range between -1.0 and 1.0, using a 0.02 min-max quantile to maintain robustness. This normalization strategy was applied during training and zero-shot evaluations for consistency. The training noise scheduler initialized from the pre-trained Stable Diffusion v2 [66] model maintained a noise strength of 0.9 and incorporated an annealed strategy to progressively reduce noise levels. We saved checkpoints every 50 iterations, with backup, validation, and visualization checkpoints set at intervals of 2,000 iterations. The training process typically converged after approximately 20,000 iterations, though we extended training to 30,000 iterations for thorough coverage. We used mean squared error (MSE) as the loss function, with reduction set to “mean” for averaged loss calculation. A customized iteration-wise exponential scheduler is applied which adjusts the learning rate iteratively using an exponential decay function. It decays the learning rate to 1% of its initial value over 25,000 iterations with a warmup phase of 100 steps. For text generation, generating a single caption for an image using LLaVA v1.6 on an RTX 3090 takes approximately 3.6 seconds, and we generate 10 captions for each image. For training, we generate 740,000 captions, using 740 GPU hours on an RTX 3090.
B.3. Evaluation metric
Following the affine-invariant depth evaluation protocol [99, 63, 90, 38, 62], for each image and the predicted relative depth $y$, we fit a pair of scalars denoting the scale and shift parameters of the transformation: $(\hat{\alpha},\hat{\beta})=g_{\psi}(y,y^{*})\in\mathbb{R}^{2}$. The metric depth prediction is obtained by $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ such that:
and mean absolute relative error, calculated as:
@@ C.2. Ablation for prompts to generate text
To study the effect of different prompts for text generation, here we use the visual question-answering model LLaVA v1.6 Mistral [49] to generate one text for each training image and testing image. The prompt we use should elicit responses that capture essential details, including the positioning of objects, their interactions, and notable features that influence monocular depth estimation. We generate different prompts using ChatGPT 4o [54], with the prompt:
@@
C.3. Different denoising steps.
As demonstrated in Figure 8, we evaluate PriorDiffusion and Marigold with different denoising steps during inference. The extracted performances are shown in Table 4. With more denoising steps, performance gradually improves. PriorDiffusion consistently outperforms Marigold across various denoising steps, converging in just 10 steps, while Marigold requires 25 steps. This suggests that the language prior could offer a good initialization, speeding up the denoising process and accelerating convergence.
Steps NYUv2 KITTI ETH3D ScanNet $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 78.7 16.9 72.2 18.0 86.0 12.7 72.1 19.2 3 92.6 8.2 77.7 15.3 92.0 8.6 87.1 11.4 4 94.2 7.0 80.4 14.1 93.6 7.6 90.7 9.3 5 94.7 6.7 82.4 13.4 94.3 7.3 91.9 8.6 10 95.1 6.4 86.5 11.9 94.9 6.9 93.2 7.8 15 95.1 6.4 87.4 11.6 95.0 6.9 93.4 7.7 20 95.2 6.3 88.0 11.4 95.0 6.9 93.7 7.5 25 95.3 6.3 88.2 11.3 95.0 6.9 93.7 7.5 50 95.3 6.3 88.5 11.2 95.0 6.9 93.8 7.5 PriorDiffusion 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 83.2 14.1 74.7 16.9 86.5 12.0 75.0 17.8 3 94.3 7.1 81.2 13.9 93.0 7.7 90.6 9.5 4 95.5 6.3 84.7 12.4 94.7 6.9 93.5 7.8 5 95.7 6.0 87.0 11.6 95.3 6.6 91.9 8.6 10 96.0 5.9 89.9 10.5 95.7 6.5 94.9 6.8 15 96.0 5.9 90.1 10.4 95.8 6.5 94.9 6.8 20 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 25 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 50 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7
Table 4: Performance for different denoising steps. PriorDiffusion consistently outperforms Marigold across various denoising steps, with a significantly faster convergence speed for the diffusion process.
C.4. Template prompt comparison.
In situations where user-provided descriptions are not available, we aim to explore whether our method can still perform effectively using either a blank input or standardized template prompts. If no specific language description that aligns with the scene is given by the user, we could instead use some pre-defined template prompt, such as “An image” or more elaborate prompts like, “A complex 3D scene with varying objects at different distances.”, as input of PriorDiffusion to maintain its performance. ” These templates, as presented in Table 5, assist in preserving the model’s performance when language input is not practical. Consequently, using template prompts as input during inference may differ from the training domain, leading to a significant covariate shift in language descriptions. This issue can be mitigated by incorporating template prompts as part of the input during training, enabling PriorDiffusion to better adapt to various scenarios where meaningful language descriptions are not feasible. Nevertheless, PriorDiffusion can achieve comparable or even better performance with Marigold when language input is not feasible, thus using a fixed prompt instead, showing that simply training with language can improve the performance of the depth estimator.
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Blank text input 95.5 6.2 89.3 10.9 95.0 6.9 “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 89.9 10.7 95.3 6.8 Template B 95.7 6.1 89.8 10.7 95.2 6.8 Template C 95.8 6.1 89.8 10.7 95.3 6.8 Marigold 95.7 6.1 89.7 10.7 95.4 6.9 With text input 96.0 5.9 90.3 10.4 95.8 6.5
  • a
  • b
  • c
Table 5: Template prompt comparison. Metrics are reported as percentages. PriorDiffusion can achieve comparable or even better performance with Marigold when language input is not feasible, thus using a fixed prompt instead. The last row is the original PriorDiffusion with text input.
C.5. Obtain language description.
To generate text descriptions for images, for training images, we use two different version of LLaVA v1.6 [49], Mistral and Vicuna, each with 5 different prompts, to generate 10 text descriptions for each training image:
  • “Describe the image in one sentence, assuming it’s a real-world image.”
  • “Provide a one-sentence description of the image, pay attention to object type, assuming it’s a real-world image.”
  • “Capture the essence of the image in a single sentence, pay attention to object relationship, assuming it’s a real-world image.”
  • “Condense the image description into one sentence, pay attention to object size, assuming it’s a real-world image.”
  • “Express the image in just one sentence, pay attention to the overall layout, assuming it’s a real-world image.”
For testing images, we use LLaVA v1.6 Mistral, then prompt this model with:
This prompt is designed to elicit responses that include essential details, such as the positioning of objects, their interactions, and notable features that may impact depth estimation. Note that all the training data we use is synthetic data, and by emphasizing “assuming it’s a real-world image,” we ensure that the descriptions align with the types of inputs and scenarios the model will encounter.